OpenStudy (anonymous):

A rectangle has an area of 102 cm2. The length of the rectangle is 17 cm. What is the perimeter of the rectangle?

OpenStudy (nirmalnema):

102=l*b 102=17*b b=120/17=6 perimeter=2(l+b) =2(17+6) 23*2 46